# One-loop QCD corrections to SSA in unweighted Drell-Yan processes

### v1

@@ Abstract
We study one-loop QCD corrections to the single transverse spin asymmetry in Drell-Yan process. The invariant mass of virtual photon and angular distributions of final lepton in Collins-Soper frame are measured. Especially, the transverse momentum of virtual photon is integrated out. Collinear twist-3 factorization formalism is adopted for the asymmetry. We use Feynman gauge in this work. To eliminate dependent twist-3 distribution functions, equation of motion for quark is used. We find that the soft divergence from the hard pole contribution in real corrections cannot be cancelled by corresponding divergences from virtual corrections. After collinear subtraction, the hard coefficient still contains soft divergence. Thus we conclude that collinear twist-3 factorization does not hold for this asymmetry at one-loop level.
@@ I Introduction
The collinear twist-3 factorization for single transverse spin asymmetry (SSA) in pion hadronic production has been proposed for many yearsEfremov and Teryaev (1982, 1985); Qiu and Sterman (1991, 1999). This factorization formalism can also be used to describe SSAs in direct photon productionQiu and Sterman (1992), semi-inclusive deeply inelastic scattering(SIDIS)Eguchi et al. (2006, 2007); Ji et al. (2006a) and Drell-Yan process(DY)Ji et al. (2006b, c) etc. However, a proof for the factorization is still missing for these processes. Different from twist-2 factorization, there are few one-loop corrections at twist-3 level in literatures for the SSA. So far, QCD corrections to several weighted SSAs in DY and SIDIS are calculated to one-loop levelVogelsang and Yuan (2009); Kang et al. (2013); Dai et al. (2015); Yoshida (2016); Chen et al. (2017, 2018). In these weighted SSAs, the transverse momentum $q_{\perp}$ for virtual photon in DY or $P_{h\perp}$ for final detected hadron in SIDIS is integrated out. It is found the twist-3 factorization really holds. These are non-trivial checks for the factorization. However, it does not imply the unweighted SSAs can also be factorized at the same order of $\alpha_{s}$. At tree level, the asymmetry for the angular distribution of final lepton in DY with $q_{\perp}$ integrated out was shown to be nonzero, please see Ma and Zhang (2015) and reference therein. In Ma and Zhang (2015), the tree level tensor structure of the hadronic tensor is made clear (see eq.(52) in the following), which has two parts: one part is proportional to the derivative of $\delta^{2}(q_{\perp})$ and the other part is proportional to $\delta^{2}(q_{\perp})$ itself. The former is called derived part and the latter is non-derivative part. Because the weight is proportional to $q_{\perp}$, the weighted SSA receives virtual correction only from derivative part. It has been illustrated in Chen et al. (2017) that this virtual correction is the same as the correction to the usual quark form factor, which is a twist-2 quantity. The direct calculation in Vogelsang and Yuan (2009) confirms this. If there is no weight, both derivative part and non-derivative part of the hadronic tensor can contribute to SSA. Thus, the examine of factorization based on weighted SSA is incomplete.
@@
In this work, we study the one-loop correction to unweighted SSA in DY with $q_{\perp}$ integrated. Explicitly, we calculate the angular distribution of final lepton. To avoid possible ambiguity for soft-gluon-pole contribution, we take Feynman gauge in this work. The method of diagram expansionQiu and Sterman (1999); Eguchi et al. (2007) is adopted here. The troubles for the expansion are mainly two aspects: one is there are many dependent twist-3 distribution functions. Some of these functions contain the bad component of quark field. The other is the gauge invariant distribution functions contain gluon field strength tensor and gauge links, but it seems impossible to recover these two quantities completely. In order to solve these two problems, we use at most one longitudinal gluon($G^{+}$) to do collinear expansion, and then use equation of motion for quark field to eliminate the bad component. After these treatments we get five independent quark-gluon-quark or quark-quark correlation functions. Three of them can be identified to $q_{\partial}$, $T_{F}$ and $T_{\Delta}$. Remaining two correlation functions are dangerous. To preserve QCD gauge invariance, the hard coefficients before these two functions must be zero. Really, our calculation confirms this. This indicates our expansion scheme preserves QCD gauge invariance. QED gauge invariance for the hadronic tensor is also checked at one-loop level. In this work, we consider the contribution proportional to $\bar{q}\otimes T_{F}$. That is, for unpolarized hadron, only the contribution from twist-2 anti-quark distribution function $\bar{q}(x)$ is considered and for polarized hadron, only twist-3 quark-gluon-quark distribution function $T_{F}$ is considered. The hard coefficients from virtual and real corrections are calculated explicitly, however, it is found after collinear subtraction the final hard coefficient still contains a divergence. Since our expansion scheme preserve both QCD and QED gauge invariance, we think the divergence indicates that collinear twist-3 factorization does not hold for the SSA we consider here. Very recently, Rein et al. (2025a, b) give the one-loop correction to a single spin asymmetry $A_{UT}$ for $E_{h}d^{3}\sigma/d^{3}P_{h}$ in lepton-hadron scattering $l+p(s_{\perp})\rightarrow h+X$. The final lepton is undetected. Please see eq.(1) of Rein et al. (2025a) for the illustration of the asymmetry. No breaking of the factorization is found there. This quantity can be studied with our method and we will study this process in near future.
@@
The structure of this paper is as follows: Sec.2 is the kinematics for lepton angular distributions in Drell-Yan process; Sec.3 is the definition of all involved twist-3 distributions and their relations resulting from equation of motion of quark field. Our expansion formalism is also presented in this section; Sec.4 contains tree level results; Sec.5 contains one-loop virtual corrections; Sec.6 contains real corrections; Sec.7 is for the renormalization of twist-3 distribution function and collinear subtraction. The final hard coefficient is also given, which contains a divergence mentioned above; In Sec.8, we give an analysis based on eikonal approximation for the hard pole contribution. It is indicated that the uncancelled divergence is a soft divergence; Sec.9 is our summary.
@@ II Kinematics
where $h_{A}$ is a spin-$\frac{1}{2}$ hadron polarized transversely with $s_{\perp}^{\mu}$ the spin vector; $h_{B}$ is a unpolarized hadron; $X$ represents undetected hadrons; the final lepton pair(here we take electron and positron as an example) is assumed from the decay of a virtual photon, and their momenta $l^{\mu}$ and $\bar{l}^{\mu}$ are detected. As usual we introduce $q^{\mu}=l^{\mu}+\bar{l}^{\mu}$ for the virtual photon. $p_{a},p_{b}$ are momenta of hadrons. The total energy squared is $s=(p_{a}+p_{b})^{2}$.
The invariant mass squared of virtual photon $Q^{2}=q^{2}$ is a hard scale. Under Bjorken limit $Q^{2}\rightarrow\infty$ and $\tau=Q^{2}/s$ fixed, we can ignore all masses of hadrons and leptonsKodaira and Tanaka (1999). The angular distribution of final lepton we want to study is
@@
$\Omega$ is the solid angle of final lepton with momentum $l^{\mu}$, defined in Collins-Soper(CS) frameCollins and Soper (1977). In this work, we take dimensional regularization to regulate ultraviolet(UV) and infrared(IR) divergences. The dimension of $q$-integration has been set to $n=4-\epsilon$. $L^{\mu\nu}$ and $W^{\mu\nu}$ are leptonic and hadronic tensors, respectively. That is,
@@
The solid angle $\Omega$ is defined in CS frame, which is a rest frame of lepton pair and is obtained from CM frame by two boostsCollins and Soper (1977). The first boost is along $Z$ axis so that $q^{z}=0$ after boost; the second boost is along the direction $\vec{q}_{\perp}$, so that $\vec{q}_{\perp}=0$ after boost. $l^{\mu}$ in CS frame is parameterized as
@@
$I\langle P_{7}\rangle=$ $\int d^{n}q\delta(q^{2}-Q^{2})P_{7,\mu\nu}W^{\mu\nu},\ P_{7}^{\mu\nu}=\frac{1}{\tilde{s}_{\perp}^{2}}\Big{(}q_{\perp}^{\mu}\tilde{s}_{\perp}^{\nu}+q_{\perp}^{\nu}\tilde{s}_{\perp}^{\mu}\Big{)},$ (10)
with $\tilde{s}_{\perp}^{\mu}=\epsilon_{\perp}^{\mu\nu}s_{\perp,\nu}$. $I\langle L\rangle$ is proportional to the differential cross section listed above, i.e.,
@@
$I\langle P_{7}\rangle$ on the other hand is proportional to the weighted cross section studied in Chen et al. (2017). Since the weighted cross section has been shown to be factorized, $I\langle P_{7}\rangle$ here is used as a check of our calculation.
@@ III Twist-3 distribution functions
$\tilde{s}_{\perp}^{\rho}T_{F}(x_{1},x_{2})=$ $g_{s}\int\frac{d\xi^{-}d\xi_{1}^{-}}{4\pi}e^{-i\xi^{-}k^{+}-i\xi_{1}^{-}k_{1}^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\mathcal{L}_{n}^{\dagger}(\xi^{-})\gamma^{+}G_{\perp}^{+\rho}(\xi^{-})\mathcal{L}_{n}(\xi^{-})\mathcal{L}^{\dagger}_{n}(\xi_{1}^{-})\psi(\xi_{1}^{-})|ps\rangle,$
$s_{\perp}^{\rho}T_{\Delta}(x_{1},x_{2})=$ $g_{s}\int\frac{d\xi^{-}d\xi_{1}^{-}}{4\pi}e^{-i\xi^{-}k^{+}-i\xi_{1}^{-}k_{1}^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\mathcal{L}_{n}^{\dagger}(\xi^{-})(-i)\gamma^{+}\gamma_{5}G_{\perp}^{+\rho}(\xi^{-})\mathcal{L}_{n}(\xi^{-})\mathcal{L}^{\dagger}_{n}(\xi_{1}^{-})\psi(\xi_{1}^{-})|ps\rangle,$ (12)
where the gauge link $\mathcal{L}_{n}$ ensures that the two distributions are gauge invariant. The definition of gauge link is
@@
with $T_{a}$ the generator of fundamental representation of $SU(N_{c})$. $P$ is path ordering operator:
@@
$-iq_{\partial}(x)\tilde{s}_{\perp}^{\rho}=$ $\int\frac{d\xi^{-}}{4\pi}e^{i\xi^{-}xp^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\gamma^{+}\gamma_{5}\partial_{\perp}^{\rho}\mathcal{L}_{n}^{\dagger}(\xi^{-})\psi(\xi^{-})|ps\rangle.$ (16)
@@
These distribution functions and relations between them can be found for example in Boer et al. (2003); Zhou et al. (2010); Kanazawa et al. (2016). $q_{T}$ and $q^{\prime}_{\partial}$ can be eliminated. Another kind of twist-3 distributions involving covariant derivative can be expressed by the distribution functions introduced aboveZhou et al. (2010). So, we expect the factorized cross section can be expressed by $T_{F},T_{\Delta}$ and $q_{\partial}$.
@@
with $\tilde{\Gamma}=\gamma_{\perp}^{\rho},\gamma_{\perp}^{\rho}\gamma_{5}$, $\Gamma=\gamma^{+},\gamma^{+}\gamma_{5}$. Then, we try to eliminate the dependent matrix elements by using equation of motion(EOM) and parity and time reversal(PT) symmetries. Such an expansion scheme can be understood because the operators in $q_{\partial}$, $T_{F}$ and $T_{\Delta}$ can always be expanded into $\bar{\psi}_{+}$, $\psi_{+}$, $G^{+}$ and $G_{\perp}$, where $\bar{\psi}_{+}$ and $\psi_{+}$ are good components of fermion field( see eq.(23)). Then, if the factorization for $q_{\partial}$, $T_{F}$ and $T_{\Delta}$ is right, preserving the expansion to a certain power of $G^{+}$ is also right, in the sense that the coefficient functions of the resulting matrix elements are finite. In our scheme, we keep the expansion to $O(G^{+})$ and $(G_{\perp})^{0}$. Details of the expansion is given below.
@@
$\int\frac{d\xi^{-}}{2\pi}e^{ik^{+}\xi^{-}}\langle Ps|\bar{\psi}_{j}(0)\psi_{i}(\xi^{-})|Ps\rangle=$ $\frac{1}{4N_{c}}\Big{[}\gamma^{-}2q(x)+\gamma_{5}\gamma_{\perp}^{\rho}s_{\perp\rho}M^{(0)}_{\gamma_{\perp}\gamma_{5}}+\gamma_{\perp}^{\rho}\tilde{s}_{\perp\rho}M^{(0)}_{\gamma_{\perp}}\Big{]}$,
$\int\frac{d\xi^{-}}{2\pi}e^{ik^{+}\xi^{-}}\langle Ps|\bar{\psi}_{j}(0)\partial_{\perp}^{\rho}\psi_{i}(\xi^{-})|Ps\rangle=$ $\frac{1}{4N_{c}}\Big{[}i\gamma_{5}\gamma^{-}s_{\perp}^{\rho}M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}+\gamma^{-}\tilde{s}_{\perp}^{\rho}M^{(0)}_{\gamma^{+},\partial_{\perp}\psi}\Big{]}$. (19)
The superscript $(0)$ implies there is no gluon. The subscript represents the gamma matrix in the correlation function and derivatives. From PT symmetry, it can be shown
$M^{(0)}_{\gamma_{\perp}}=M^{(0)}_{\gamma^{+},\partial\psi}=0.$ (20)
@@
$g_{s}\int\frac{d\xi^{-}}{2\pi}\int\frac{d\xi_{1}^{-}}{2\pi}e^{ik^{+}\xi^{-}+ik_{1}^{+}\xi_{1}^{-}}\langle ps|\bar{\psi}_{j}(0)G_{a}^{+}(\xi^{-})\psi_{i}(\xi_{1}^{-})|ps\rangle=\frac{T_{a}}{4N_{c}C_{F}}\Big{[}i\gamma_{\perp}^{\rho}\tilde{s}_{\perp\rho}M^{(1)}_{\gamma_{\perp}}+\gamma_{5}\gamma_{\perp}^{\rho}s_{\perp\rho}M^{(1)}_{\gamma_{\perp}\gamma_{5}}\Big{]}$; (21)
@@
$g_{s}\int\frac{d\xi^{-}}{2\pi}\int\frac{d\xi_{1}^{-}}{2\pi}e^{ik^{+}\xi^{-}+ik_{1}^{+}\xi_{1}^{-}}\langle ps|\bar{\psi}_{j}(0)G_{a}^{+}(\xi^{-})\partial_{\perp}^{\rho}\psi_{i}(\xi_{1}^{-})|ps\rangle=$ $\frac{T_{a}}{4N_{c}C_{F}}\Big{[}\gamma^{-}\tilde{s}_{\perp}^{\rho}M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}+i\gamma_{5}\gamma^{-}s_{\perp}^{\rho}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}\Big{]}$;
$g_{s}\int\frac{d\xi^{-}}{2\pi}\int\frac{d\xi_{1}^{-}}{2\pi}e^{ik^{+}\xi^{-}+ik_{1}^{+}\xi_{1}^{-}}\langle ps|\bar{\psi}_{j}(0)[\partial_{\perp}^{\rho}G_{a}^{+}(\xi^{-})]\psi_{i}(\xi_{1}^{-})|ps\rangle=$ $\frac{T_{a}}{4N_{c}C_{F}}\Big{[}\gamma^{-}\tilde{s}_{\perp}^{\rho}M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}+i\gamma_{5}\gamma^{-}s_{\perp}^{\rho}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}\Big{]}$. (22)
From PT symmetry of QCD, all of these $M^{(i)}$ are real. Note that for $M^{(1)}$, one $g_{s}$ is included in the matrix element.
@@
The $\doteq$ implies $G_{\perp}$ and higher order of $G^{+}$ such as $(G^{+})^{2}$ terms are ignored on right hand side. P means principal value, and $k_{2}^{+}=k^{+}+k_{1}^{+}$. To one-loop level, a consistent treatment of $\gamma_{5}$ is important. We adopt HVBM scheme’t Hooft and Veltman (1972); Breitenlohner and Maison (1977) in this work. In this scheme $\gamma_{5}$ is defined as a four dimensional quantity. In addition, spin vectors $s_{\perp}^{\mu}$ and $\tilde{s}_{\perp}^{\mu}$ are also defined as four dimensional quantities. Thus following identities can be applied
@@
which reduce the number of gamma matrices.

### v2

@@ Abstract
We study one-loop QCD corrections to the single transverse spin asymmetry in Drell-Yan process. The invariant mass of virtual photon and angular distributions of final lepton in Collins-Soper frame are measured. Especially, the transverse momentum of virtual photon is integrated out. Collinear twist-3 factorization formalism is adopted for the asymmetry. We use Feynman gauge in this work. To eliminate dependent twist-3 distribution functions, equation of motion for quark is used. The results satisfy both QED and QCD gauge invariance. It is confirmed that all divergences from virtual and real corrections can be removed consistently by collinear subtraction. Finite hard coefficients for the convolution of $\bar{q}(x)$ and $T_{F}(x_{1},x_{2})$ are presented.
@@ I Introduction
The collinear twist-3 factorization for single transverse spin asymmetry (SSA) in pion hadronic production has been proposed for many years[1, 2, 3, 4]. This factorization formalism can also be used to describe SSAs in direct photon production[5], semi-inclusive deeply inelastic scattering(SIDIS)[6, 7, 8] and Drell-Yan process(DY)[9, 10] etc. However, a proof for the factorization is still missing for these processes. Different from twist-2 factorization, there are few one-loop corrections at twist-3 level in literatures for the SSA. So far, QCD corrections to several weighted SSAs in DY and SIDIS are calculated to one-loop level[11, 12, 13, 14, 15, 16]. In these weighted SSAs, the transverse momentum $q_{\perp}$ for virtual photon in DY or $P_{h\perp}$ for final detected hadron in SIDIS is integrated out. It is found the twist-3 factorization really holds. These are non-trivial checks for the factorization. However, it does not imply the unweighted SSAs can also be factorized at the same order of $\alpha_{s}$. At tree level, the asymmetry for the angular distribution of final lepton in DY with $q_{\perp}$ integrated out was shown to be nonzero, please see [17] and reference therein. In [17], the tree level tensor structure of the hadronic tensor is made clear (see eq.(55) in the following), which has two parts: one part is proportional to the derivative of $\delta^{2}(q_{\perp})$ and the other part is proportional to $\delta^{2}(q_{\perp})$ itself. The former is called derived part and the latter is non-derivative part. Because the weight is proportional to $q_{\perp}$, the weighted SSA receives virtual correction only from derivative part. It has been illustrated in [15] that this virtual correction is the same as the correction to the usual quark form factor, which is a twist-2 quantity. The direct calculation in [11] confirms this. If there is no weight, both derivative part and non-derivative part of the hadronic tensor can contribute to SSA. Thus, the examine of factorization based on weighted SSA is incomplete. Moreover, unweighted SSA is expected to have smaller errors than weighted SSA in experiments. One can consult [18] for recent experimental results for SSA in DY and [19] for preliminary weighted SSA.
@@
In this work, we study the one-loop correction to unweighted SSA in DY with $q_{\perp}$ integrated. Explicitly, we calculate the angular distribution of final lepton. To avoid possible ambiguity for soft-gluon-pole contribution, we take Feynman gauge in this work. The method of diagram expansion[4, 7] is adopted here. The troubles for the expansion are mainly two aspects: one is there are many dependent twist-3 distribution functions. Some of these functions contain the bad component of quark field. The other is the gauge invariant distribution functions contain gluon field strength tensor and gauge links, but it seems impossible to recover these two quantities completely. In order to solve these two problems, we use at most one longitudinal gluon($G^{+}$) to do collinear expansion, and then use equation of motion for quark field to eliminate the bad component. After these treatments we get five independent quark-gluon-quark or quark-quark correlation functions. Three of them can be identified to $q_{\partial}$, $T_{F}$ and $T_{\Delta}$, which are gauge invariant. Remaining two correlation functions are not gauge invariant and dangerous to the factorization assumption. To preserve QCD gauge invariance, the hard coefficients before these two functions must be zero. Really, our calculation confirms this. This indicates our expansion scheme preserves QCD gauge invariance. QED gauge invariance for the hadronic tensor is also checked at one-loop level, i.e., $q^{\mu}W_{\mu\nu}=0$, with $W^{\mu\nu}$ the hadronic tensor. In this work, we consider the contribution proportional to $\bar{q}\otimes T_{F}$. That is, for unpolarized hadron, only the contribution from twist-2 anti-quark distribution function $\bar{q}(x)$ is considered and for polarized hadron, only twist-3 quark-gluon-quark distribution function $T_{F}$ is considered. The hard coefficients from virtual and real corrections are calculated explicitly. After collinear subtraction, all divergences from virtual and real corrections are removed consistently.
Very recently, [20, 21] give the one-loop correction to a single spin asymmetry $A_{UT}$ for $E_{h}d^{3}\sigma/d^{3}P_{h}$ in lepton-hadron scattering $l+p(s_{\perp})\rightarrow h+X$. The final lepton is undetected. Please see eq.(1) of [20] for the illustration of the asymmetry. Our method for collinear expansion is different from theirs. In [20, 21], light-cone gauge is adopted. But we really find many common features for the one-loop corrections. For example, there are nonpole or integral contributions in virtual corrections, which come from the absorptive part of loop integrals rather than from an on-shell propagator. Especially, these nonpole contributions contain collinear divergences and have a different tensor structure from tree level hadronic tensor. However, these divergences can still be subtracted by one-loop corrections to $T_{F}(x,x)$. This is a new feature of twist-3 calculation. It is important for us to extend the method of this work to lepton-hadron scattering in future.
@@
The structure of this paper is as follows: Sec.2 is the kinematics for lepton angular distributions in Drell-Yan process; Sec.3 is the definition of all involved twist-3 distributions and their relations resulting from equation of motion of quark field. Our expansion formalism is also presented in this section; Sec.4 contains tree level results; Sec.5 contains one-loop virtual corrections; Sec.6 contains real corrections; Sec.7 is for the renormalization of twist-3 distribution function and collinear subtraction. The final finite hard coefficients are also given; Sec.8 is our summary.
@@ II Kinematics
where $h_{A}$ is a spin-$\frac{1}{2}$ hadron polarized transversely with $s_{\perp}^{\mu}$ the spin vector; $h_{B}$ is an unpolarized hadron; $X$ represents undetected hadrons; the final lepton pair(here we take electron and positron as an example) is assumed from the decay of a virtual photon, and their momenta $l^{\mu}$ and $\bar{l}^{\mu}$ are detected. As usual we introduce $q^{\mu}=l^{\mu}+\bar{l}^{\mu}$ for the virtual photon. $p_{a},p_{b}$ are momenta of hadrons. The total energy squared is $s=(p_{a}+p_{b})^{2}$.
The invariant mass squared of virtual photon $Q^{2}=q^{2}$ is a hard scale. Under Bjorken limit $Q^{2}\rightarrow\infty$ and $\tau=Q^{2}/s$ fixed, we can ignore all masses of hadrons and leptons[22]. The angular distribution of final lepton we want to study is
@@
$\Omega$ is the solid angle of final lepton with momentum $l^{\mu}$, defined in Collins-Soper(CS) frame[23]. In this work, we take dimensional regularization to regulate ultraviolet(UV) and infrared(IR) divergences. The dimension of $q$-integration has been set to $n=4-\epsilon$. $L^{\mu\nu}$ and $W^{\mu\nu}$ are leptonic and hadronic tensors, respectively. That is,
@@
The solid angle $\Omega$ is defined in CS frame, which is a rest frame of lepton pair and is obtained from CM frame by two boosts[23]. The first boost is along $Z$ axis so that $q^{z}=0$ after boost; the second boost is along the direction $\vec{q}_{\perp}$, so that $\vec{q}_{\perp}=0$ after boost. $l^{\mu}$ in CS frame is parameterized as
@@
$I\langle P_{7}\rangle=$ $\int d^{n}q\delta(q^{2}-Q^{2})P_{7,\mu\nu}W^{\mu\nu},\ P_{7}^{\mu\nu}=\frac{1}{\tilde{s}_{\perp}^{2}}\Big(q_{\perp}^{\mu}\tilde{s}_{\perp}^{\nu}+q_{\perp}^{\nu}\tilde{s}_{\perp}^{\mu}\Big).$ (10)
$I\langle L\rangle$ is proportional to the differential cross section listed above, i.e.,
@@
$I\langle P_{7}\rangle$ on the other hand is proportional to the weighted cross section studied in [15]. Since the weighted cross section has been shown to be factorized, $I\langle P_{7}\rangle$ here is used as a check of our calculation.
@@ III Twist-3 distribution functions
$\tilde{s}_{\perp}^{\rho}T_{F}(x_{1},x_{2})=$ $g_{s}\int\frac{d\xi^{-}d\xi_{1}^{-}}{4\pi}e^{i\xi^{-}k^{+}+i\xi_{1}^{-}k_{1}^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\mathcal{L}_{n}^{\dagger}(\xi^{-})\gamma^{+}G_{\perp}^{+\rho}(\xi^{-})\mathcal{L}_{n}(\xi^{-})\mathcal{L}^{\dagger}_{n}(\xi_{1}^{-})\psi(\xi_{1}^{-})|ps\rangle,$
$s_{\perp}^{\rho}T_{\Delta}(x_{1},x_{2})=$ $g_{s}\int\frac{d\xi^{-}d\xi_{1}^{-}}{4\pi}e^{i\xi^{-}k^{+}+i\xi_{1}^{-}k_{1}^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\mathcal{L}_{n}^{\dagger}(\xi^{-})(-i)\gamma^{+}\gamma_{5}G_{\perp}^{+\rho}(\xi^{-})\mathcal{L}_{n}(\xi^{-})\mathcal{L}^{\dagger}_{n}(\xi_{1}^{-})\psi(\xi_{1}^{-})|ps\rangle,$ (12)
where $G_{\perp}^{+\rho}=G_{a\perp}^{+\rho}T_{a}$ is gluon field strength tensor and $\mathcal{L}_{n}$ is the gauge link which ensures that the two distributions are gauge invariant. The definition of gauge link is
@@
with $G_{a}^{\mu}$ gluon field and $T_{a}$ the generator of fundamental representation of $SU(N_{c})$. $P$ is path ordering operator:
@@
$-iq_{\partial}(x)s_{\perp}^{\rho}=$ $\int\frac{d\xi^{-}}{4\pi}e^{i\xi^{-}xp^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\gamma^{+}\gamma_{5}\partial_{\perp}^{\rho}\mathcal{L}_{n}^{\dagger}(\xi^{-})\psi(\xi^{-})|ps\rangle.$ (16)
@@
These distribution functions and relations between them can be found for example in [24, 25, 26]. $q_{T}$ and $q^{\prime}_{\partial}$ can be eliminated. Another kind of twist-3 distributions involving covariant derivative can be expressed by the distribution functions introduced above. So, we expect the factorized cross section can be expressed by $T_{F},T_{\Delta}$ and $q_{\partial}$. In [26], $T_{F},T_{\Delta}$ are called dynamical twist-3 distribution functions, while $q_{\partial},q^{\prime}_{\partial}$ and $q_{T}$ are kinematical ones. With the notation of [26], our twist-3 distribution functions are expressed as
$T_{F}(x_{1},x_{2})=-2\pi MF_{FT}^{q}(x_{2},x_{1}),\ T_{\Delta}(x_{1},x_{2})=2\pi MG_{FT}^{q}(x_{2},x_{2}),$
$q_{T}(x)=Mg_{T}(x),\ q_{\partial}(x)=Mg_{1T}^{(1),q}(x),\ q^{\prime}_{\partial}(x)=Mf_{1T}^{\perp(1),q}(x),$ (18)
where $M$ is hadron mass. The minus sign for $T_{F}$ relation is from different definition of covariant derivatives. We have $D^{\mu}\equiv\partial^{\mu}+ig_{s}G^{\mu}$, but [26] has $D^{\mu}\equiv\partial^{\mu}-ig_{s}G^{\mu}$. That is, the two strong couplings are opposite. The relation for $q^{\prime}_{\partial}(x)$ holds if the gauge links in $q^{\prime}_{\partial}$ and in Sivers function $f_{1T}^{\perp,q}(x,k_{\perp})$ point to -$\infty^{-}$ or $\infty^{-}$ at the same time.
@@
with $\tilde{\Gamma}=\gamma_{\perp}^{\rho},\gamma_{\perp}^{\rho}\gamma_{5}$, $\Gamma=\gamma^{+},\gamma^{+}\gamma_{5}$. Then, we try to eliminate the dependent matrix elements by using equation of motion(EOM) and parity and time reversal(PT) symmetries. Such an expansion scheme can be understood because the operators in $q_{\partial}$, $T_{F}$ and $T_{\Delta}$ can always be expanded into $\bar{\psi}_{+}$, $\psi_{+}$, $G^{+}$ and $G_{\perp}$, where $\bar{\psi}_{+}$ and $\psi_{+}$ are good components of fermion field( see eq.(24)). Then, if the factorization for $q_{\partial}$, $T_{F}$ and $T_{\Delta}$ is right, preserving the expansion to a certain power of $G^{+}$ is also right, in the sense that the coefficient functions of the resulting matrix elements are finite. In our scheme, we keep the expansion to $O(G^{+})$ and $O((G_{\perp})^{0})$. Details of the expansion is given below.
@@
$\int\frac{d\xi^{-}}{2\pi}e^{ik^{+}\xi^{-}}\langle ps|\bar{\psi}_{j}(0)\psi_{i}(\xi^{-})|ps\rangle=$ $\frac{\delta_{ij}}{4N_{c}}\Big[\gamma^{-}2q(x)+\gamma_{5}\gamma_{\perp}^{\rho}s_{\perp\rho}M^{(0)}_{\gamma_{\perp}\gamma_{5}}(k^{+})+\gamma_{\perp}^{\rho}\tilde{s}_{\perp\rho}M^{(0)}_{\gamma_{\perp}}(k^{+})\Big]_{ij},$
$\int\frac{d\xi^{-}}{2\pi}e^{ik^{+}\xi^{-}}\langle ps|\bar{\psi}_{j}(0)\partial_{\perp}^{\rho}\psi_{i}(\xi^{-})|ps\rangle=$ $\frac{\delta_{ij}}{4N_{c}}\Big[i\gamma_{5}\gamma^{-}s_{\perp}^{\rho}M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+})+\gamma^{-}\tilde{s}_{\perp}^{\rho}M^{(0)}_{\gamma^{+},\partial_{\perp}\psi}(k^{+})\Big]_{ij},$ (20)
with $k^{+}=xp^{+}$. The superscript $(0)$ implies there is no gluon. The subscript represents the gamma matrix in the correlation function, and the fields with transverse derivative are also shown. From PT symmetry, it can be shown
$M^{(0)}_{\gamma_{\perp}}(k^{+})=M^{(0)}_{\gamma^{+},\partial_{\perp}\psi}(k^{+})=0.$ (21)
@@
$g_{s}\int\frac{d\xi^{-}d\xi_{1}^{-}}{(2\pi)^{2}}e^{ik^{+}\xi^{-}+ik_{1}^{+}\xi_{1}^{-}}\langle ps|\bar{\psi}_{j}(0)G_{a}^{+}(\xi^{-})\psi_{i}(\xi_{1}^{-})|ps\rangle=\frac{T^{a}_{ij}}{4N_{c}C_{F}}\Big[i\gamma_{\perp}^{\rho}\tilde{s}_{\perp\rho}M^{(1)}_{\gamma_{\perp}}(k^{+},k_{1}^{+})+\gamma_{5}\gamma_{\perp}^{\rho}s_{\perp\rho}M^{(1)}_{\gamma_{\perp}\gamma_{5}}(k^{+},k_{1}^{+})\Big]_{ij};$ (22)
@@
$g_{s}\int\frac{d\xi^{-}d\xi_{1}^{-}}{(2\pi)^{2}}e^{ik^{+}\xi^{-}+ik_{1}^{+}\xi_{1}^{-}}\langle ps|\bar{\psi}_{j}(0)G_{a}^{+}(\xi^{-})\partial_{\perp}^{\rho}\psi_{i}(\xi_{1}^{-})|ps\rangle$
$=\frac{T^{a}_{ij}}{4N_{c}C_{F}}\Big[\gamma^{-}\tilde{s}_{\perp}^{\rho}M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+i\gamma_{5}\gamma^{-}s_{\perp}^{\rho}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})\Big]_{ij};$
$g_{s}\int\frac{d\xi^{-}d\xi_{1}^{-}}{(2\pi)^{2}}e^{ik^{+}\xi^{-}+ik_{1}^{+}\xi_{1}^{-}}\langle ps|\bar{\psi}_{j}(0)[\partial_{\perp}^{\rho}G_{a}^{+}(\xi^{-})]\psi_{i}(\xi_{1}^{-})|ps\rangle$
$=\frac{T^{a}_{ij}}{4N_{c}C_{F}}\Big[\gamma^{-}\tilde{s}_{\perp}^{\rho}M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})+i\gamma_{5}\gamma^{-}s_{\perp}^{\rho}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big]_{ij}.$ (23)
From PT symmetry of QCD, all $M^{(i)}$ are real. Note that all $M^{(1)}$ are proportional to $g_{s}$.
@@
The $\doteq$ implies $G_{\perp}$ and higher order of $G^{+}$ such as $(G^{+})^{2}$ terms are ignored on right hand side. P means principal value(PV), and $k_{2}^{+}=k^{+}+k_{1}^{+}$. To one-loop level, a consistent treatment of $\gamma_{5}$ is important. We adopt HVBM scheme[27, 28] in this work. In this scheme $\gamma_{5}$ is defined as a four dimensional quantity. In addition, spin vectors $s_{\perp}^{\mu}$ and $\tilde{s}_{\perp}^{\mu}$ are also defined as four dimensional quantities. Thus following identities can be applied
@@
which reduce the number of gamma matrices. Our conventions for $\gamma_{5}$ are listed in Appendix.C.